Use the rational exponent equation L=2.43*W^0.3326 to estimate the wingspan L in feet of the bird that weighs W pounds. My bird weighs 2 1/3 ozs. L=2.43*2.33^0.3326
L=3.22 What? It is supposed to be feet. However, a robin, a bird that weighs 2 1/3 ozs, can not have a wing span of 3.22 feet? What am I doing wrong?
Please help?

Answer by stanbon(48535)

Change 2 1/3 ounces to pounds and try again.
2 1/3 = 7/3 ounces
(7/3) oz / 16 oz = (7/48) lb.
Put that in your formula.
Cheers,
Stan H.